# Link Prediction with Physics-Inspired Graph Neural Networks thanks:

### v2

@@ V-B Results
More formally, by minimizing (4), we expect the squared norm of the positive (negative) gradients to decrease (increase). To measure this we take advantage of the AUROC metric, considering the positive as class 0, while the negatives as class 1, through the AUROC we evaluate how much these two classes are separated based on their scores, and regardless of any threshold. In particular if we consider as ground truth a stack of 1’s and 0’s for negatives and positives $\mathcal{S}_{\mathcal{E}_{t}}=\{\mathbb{1}_{\{(i,j)\in\mathcal{E}_{neg}\}}|\forall(i,j)\in\mathcal{E}_{t}\}$, $GS$ at time $t$ is computed as
@@
$GS^{t}_{ht,hm}$ measures the ability to classify as 0 the positive heterophilic edges, and as 1 the negative homophilic edges. In other words, it measures how much we can separate the positive and the negative edges based on the squared norm of their edge gradient. We provide a visual example in Figure 2, we consider 1 on the y-axis as the homophilic edges, while the 0 y-axis refers to the heterophilic edges. We visualize a total of 50 samples, but the scores refers to those computed within the whole test set. These are produced from a fully-trained GRAFF-LP, monitoring $GS^{t}_{hm,hm}$ and $GS^{t}_{ht,ht}$. Figure 3, also includes the distributions of the squared norm gradients, to better understand how they separate.
@@
Figure 2: $||(\nabla\mathbf{H}^{t})_{i,j}||^{2}$ evolution with a fully-trained 9-layers GRAFF-LP via $f_{g}$ on Minesweeper.
@@
These visualizations, along with those in Appendix IV-C, provide qualitative and quantitative evidences that GRAFF-LP induces edge attraction and repulsion via $f_{g}$ as the nodes go through the network’s layers . We answer positively to RQ2.
@@
In Table IV, instead we show $GS^{T}$, which is the total gradient separability computed after the last message-passing layer, for each model and averaged across 10 random seeds. We report the results and comments about Tolokers in Appendix IV-A. In this analysis, we do not use any statistical test, since we do not want to assess what model have the highest $GS^{T}$, we want to understand what models have a sufficiently high value of $GS^{T}$ to acknowledge their ability to separate edge gradients. We consider $GS^{T}>90\%$ as an evidence for the model to be able of distinguishing the nature of the edges, based uniquely on the squared norm of the gradients. The most interesting aspect of $GS^{T}$, is the consistent increase that we record for GRAFF-LP, which we associate with the capability of this model to understand the negative and positive edges based on the evolution of their gradients. Indeed, obtaining $GS^{T}=100\%$ implies that the edges can be classified as positive or not, based solely on $\{\nabla^{T}_{pos},\nabla^{T}_{neg}\}$, which are available even before the Decoding phase. According to our results, GRAFF-LP allows to do so even though is not trained explicitly to do it. Equation (4) lets the model to minimize the squared norm of the gradients, but it refers to the message-passing edges, not to the unseen edges that we want to predict. In the Roman Empire experiments, when we go from $f_{h}$ to $f_{g}$ the AUROC improvement is limited to $+1.02\%$ (see Table III), but when it comes to gauging the gradient separability difference, we have a $+65.52\%$. Since the only difference in the two models is the readout, we affirm that $f_{g}$ induces the model to learn to separate the edges following a strategy driven by the gradients. The only dataset where GRAFF-LP seems to not catch this behavior is Questions, where $GS^{T}$ is close to 0. Anyway, this is not an issue, since the AUROC is faulty when it approaches 50%, in our case it means that the negative edges decrease, rather then the positive ones. For this reason, we conjecture that GRAFF-LP is learning the opposite behavior. The same happens to GCN, which is even closer to 0 than GRAFF-LP since it reaches down to $GS^{T}=14.32\%$. Surprisingly, not only our model is able to learn to separate the edge gradients, but also the other node-based methods, except for the MLP, ELPH, NCNC. Nonetheless, NCNC scores $GS^{T}>90$ in Minesweeper. Despite this, we recognize GRAFF-LP as the approach that consistently show this behavior, we can conclude that this is due to the PIrd message-passing and readout schemes. ELPH, reaches state-of-the-art performance in Minesweeper, leaving GRAFF-LP as the second best performing model, but our model offers more transparency in the model’s behavior, as confirmed by $GS^{T}$. These results, describe the full contribution of our newly introduced readout function $f_{g}$ answering to RQ1. Specifically, $f_{g}$ can enhance the prediction performance (see Table III), and also the model’s transparency at the inference phase, as illustrated in this Section.
@@
TABLE IV: Comparison of $GS^{T}$, when the model is trained with $f_{h}$ or $f_{g}$. The score that is the closest to 100 or 0 is highlighted in bold. We report the percentage variation between the results of $f_{h}$ and $f_{g}$ as $\Delta$.
@@
Runtime Analysis
We briefly show some results about time and space complexities that these baselines have. We already discussed space complexity in terms of parameters, but in this Section, we report the real-world values associated with the experiments. In Table V, we show the number of parameters of each model and the inference time, computed on the test edges for 10 epochs. All the models have $L=3$, $L_{MLP}=1$, and $d_{h}=d_{MLP}=64$, to make a fair comparison across models. What we notice is that according to our discussion on complexity, GRAFF-LP is the lightest model, and also comparable with the other node-based methods such as GCN, GAT and GraphSAGE. ELPH, which is a subgraph-based method has the highest inference time, as expected. While NCNC, even though it is known to be more efficient than ELPH, and more similar to the node-based methods, it still presents a significant gap in terms of inference, due to the structural features computation. Furthermore, it leads to out-of-memory issues in Tolokers. In Appendix V, we report additional results on runtime analysis.
@@
Does Class Heterophily Impact on GNN performance?
In the premises of this paper, we questioned on the absence of explicit methods that tackle heterophily in link prediction. This is a gap in the literature that is worth studying since we have seen in the previous Sections, that in some datasets, even subgraph-based approaches fail to achieve competitive performance. Other than the mere performance, in the literature is pointed out the complicated nature of the task [8], namely, how can i understand the latent factors that connect two entities? In this paper, we focused on datasets, that are known to be heterophilic for node classification. Are they representative enough for the link prediction under a heterophily scenario even if $\xi_{adj}$ is determined only from class labels? Does simple GNN baselines struggle to learn from these graphs, in the same fashion as node classification? In other words, can they learn to predict existing(non-existing) links when two entities have different(same) classes? These concerns are related to RQ3. To answer it, we measure the AUROC that we obtain when we try to classify all the possible class mixes, namely when the positive or negative edges are homophilic or heterophilic. In this way, it is possible to measure 4 different AUROCs that are representative of the model’s ability to separate edges of different natures.
Let us define $AUC_{\mathcal{U},\mathcal{V}}$, as the ability of a binary classifier to correctly distinguish the sets $\mathcal{U}$ and $\mathcal{V}$ as existing or non-existing edges. Here the ground truth is that $\mathcal{U}$ contains positives and $\mathcal{V}$ the negatives. The nature of these can be homophilic or heterophilic, namely $\mathcal{U},\mathcal{V}\in\{hm,ht\}$. In Figure 4, we take into consideration Roman Empire as an example of a dataset where the models fail to achieve high performance, and we show that is not due to heterophilic edges (i.e. $AUC_{ht,hm},AUC_{ht,ht}$), since we have poor performance also on the homophilic ones $AUC_{hm,ht},AUC_{hm,hm}$. In Appendix IV-D, we provide additional evidence of this behavior, that it presents also when models reach high performance. Generally, we understand that the performance do not vary significantly, and the models’ rankings remain always consistent. We conclude that models that achieve competitive performances in homophilic link prediction do not perform as well in our benchmark, but this is not due to class heterophily. This result answers negatively to RQ3 and highlights the need for new homophilic measures that can better identify datasets representative of link prediction under heterophily, because of its relevance in several applications.
@@ VI Conclusions
This paper presented GRAFF-LP, a link prediction model built upon GNNs and driven by a PIrd bias. Our method is designed to handle heterophilic graphs, enriching the current literature available on the topic, which is often confined to node classification. We show that GRAFF-LP outperforms the other baselines in 3 out of 4 datasets. We also propose $f_{g}$, a PIrd readout function based on the edge gradients, that without any tuning is able to improve the GRAFF-LP performance. Surprisingly, we observe that other models also benefit from $f_{g}$, but GRAFF-LP consistently ranks among the top 2. To better understand the effect of $f_{g}$, we study the evolution of the squared norm of the edge gradients via a novel metric named $GS$, and we surprisingly found that, under the lens of this norm, the GRAFF-LP’s message-passing separate the positive and negatives edges. GRAFF-LP showcases this behavior more consistently than the other baselines, favored by the attraction and repulsion bias of the gradient flow interpretation. We also brought some evidence that the baselines do not struggle to perform link prediction because of class heterophily. This result exposes the need for new measures that help identify challenging scenarios to advance our knowledge of link prediction under heterophily. Future works should address the aspects related to new heterophily metrics based on the input features since they are not task-dependent. Moreover, we plan to further understand how GRAFF-LP is able to learn the edge gradient separation, even though it was not trained explicitly to do so. Then, it would be interesting to adapt other PIrd GNNs or develop new ones specifically suited for link prediction.

### v3

@@ V-B2 Physics-Inspired Link Prediction
More formally, by minimizing (4), we expect the squared norm of the positive (negative) gradients to decrease (increase). To measure this, we take advantage of the AUROC metric, considering the positives as class 0, while the negatives as class 1. Through the AUROC, we evaluate how much these two classes are separated based on their scores, regardless of any threshold. In particular if we consider as ground truth a stack of 1’s and 0’s for negatives and positives $\mathcal{S}_{\mathcal{E}_{t}}=\{\mathbb{1}_{\{(i,j)\in\mathcal{E}_{neg}\}}|\forall(i,j)\in\mathcal{E}_{t}\}$, $GS$ at time $t$ is computed as
@@
$GS^{t}_{ht,hm}$ measures the ability to classify as 0 the positive heterophilic edges, and as 1 the negative homophilic edges. In other words, it measures how much we can separate the positive and the negative edges based on the squared norm of their edge gradient. We provide a visual example in Figure 2, where we consider 1 on the y-axis as the homophilic edges, while the 0 y-axis refers to the heterophilic edges. We visualize a total of 50 samples, but the scores refer to those computed within the whole test set. These are produced from a fully-trained GRAFF-LP, monitoring $GS^{t}_{hm,hm}$ and $GS^{t}_{ht,ht}$. Figure 3 also includes the distributions of the squared norm gradients to better understand how they separate.
@@
Figure 2: $||(\nabla\mathbf{H}^{t})_{i,j}||^{2}$ evolution with a fully-trained 9-layer GRAFF-LP via $f_{g}$ on Minesweeper.
@@
These visualizations, along with those in Appendix -E, provide qualitative and quantitative evidence that GRAFF-LP induces edge attraction and repulsion via $f_{g}$ as the nodes go through the network’s layers. We answer positively to RQ2.
@@
In Table IV, instead we show $GS^{T}$, which is the total gradient separability computed after the last message-passing layer, for each model and averaged across 10 random seeds. We report the results and comments about Tolokers in Appendix -C. In this analysis, we do not use any statistical test, since we do not want to assess which model has the highest $GS^{T}$; we want to understand what models have a sufficiently high value of $GS^{T}$ to acknowledge their ability to separate edge gradients. We consider $GS^{T}>90\%$ as evidence for the model to be able to distinguish the nature of the edges, based uniquely on the squared norm of the gradients. The most interesting aspect of $GS^{T}$ is the consistent increase that we record for GRAFF-LP, which we associate with the capability of this model to understand the negative and positive edges based on the evolution of their gradients. Indeed, obtaining $GS^{T}=100\%$ implies that the edges can be classified as positive or not, based solely on $\{\nabla^{T}_{pos},\nabla^{T}_{neg}\}$, which are available even before the Decoding phase. According to our results, GRAFF-LP allows us to do so even though it is not trained explicitly to do it. Equation (4) lets the model minimize the squared norm of the gradients, but it refers to the message-passing edges, not to the unseen edges that we want to predict. In the Roman Empire experiments, when we go from $f_{h}$ to $f_{g}$ the AUROC improvement is limited to $+1.02\%$ (see Table III), but when it comes to gauging the gradient separability difference, we have a $+65.52\%$. Since the only difference in the two models is the readout, we affirm that $f_{g}$ induces the model to learn to separate the edges following a strategy driven by the gradients. The only dataset where GRAFF-LP seems not to catch this behavior is Questions, where $GS^{T}$ is close to 0. Anyway, this is not an issue, since the AUROC is faulty when it approaches 50%, in our case, it means that the negative edges decrease, rather than the positive ones. For this reason, we conjecture that GRAFF-LP is learning the opposite behavior. The same happens to GCN, which is even closer to 0 than GRAFF-LP since it reaches down to $GS^{T}=14.32\%$. Surprisingly, not only is our model able to learn to separate the edge gradients, but also the other node-based methods, except for the MLP, ELPH, NCNC. Nonetheless, NCNC scores $GS^{T}>90$ in Minesweeper. Despite this, we recognize GRAFF-LP as the approach that consistently shows this behavior; we can conclude that this is due to the PIrd message-passing and readout schemes. ELPH reaches state-of-the-art performance in Minesweeper, leaving GRAFF-LP as the second best performing model, but our model offers more transparency in the model’s behavior, as confirmed by $GS^{T}$. These results describe the full contribution of our newly introduced readout function $f_{g}$ answering RQ1. Specifically, $f_{g}$ can enhance the prediction performance (see Table III), and also the model’s transparency at the inference phase, as illustrated in this Section.
@@
TABLE IV: Comparison of $GS^{T}$, when the model is trained with $f_{h}$ or $f_{g}$. The score that is closest to 100 or 0 is highlighted in bold. We report the percentage variation between the results of $f_{h}$ and $f_{g}$ as $\Delta$.
@@
V-B3 Runtime Analysis
We briefly show some results about the time and space complexities that these baselines have. We already discussed space complexity in terms of parameters, but in this Section, we report the real-world values associated with the experiments. In Table V, we show the number of parameters of each model and the inference time, computed on the test edges for 10 epochs. All the models have $L=3$, $L_{MLP}=1$, and $d_{h}=d_{MLP}=64$, to make a fair comparison across models. What we notice is that according to our discussion on complexity, GRAFF-LP is the lightest model, and also comparable with the other node-based methods such as GCN, GAT and GraphSAGE. ELPH, which is a subgraph-based method has the highest inference time, as expected. While NCNC, even though it is known to be more efficient than ELPH, and more similar to the node-based methods, it still presents a significant gap in terms of inference, due to the computation of structural features. Furthermore, it leads to out-of-memory issues in Tolokers. In Appendix Extended Runtime Analysis, we report additional results on runtime analysis.
@@ V-B3 Runtime Analysis
V-B4 Does Class Heterophily Impact GNN performance?
In this paper’s premises, we questioned the absence of explicit methods that tackle heterophily in link prediction. This is a gap in the literature that is worth studying since we have seen in the previous Sections, that in some datasets, even subgraph-based approaches fail to achieve competitive performance. Other than the mere performance, in the literature is pointed out the complicated nature of the task [6], namely, how can the latent factors that connect two entities be understood? In this paper, we focused on datasets that are known to be heterophilic for node classification. Are they representative enough for link prediction under heterophily, even if $\xi_{adj}$ is determined only from class labels? Do simple GNN baselines struggle to learn from these graphs, in the same fashion as node classification? In other words, can they learn to predict existing(non-existing) links when two entities have different(same) classes? These concerns are related to RQ3. To answer it, we measure the AUROC that we obtain when we try to classify all the possible class mixes, namely when the positive or negative edges are homophilic or heterophilic. In this way, it is possible to measure 4 different AUROCs that are representative of the model’s ability to separate edges of different natures.
Let us define $AUC_{\mathcal{U},\mathcal{V}}$ as the ability of a binary classifier to correctly distinguish the sets $\mathcal{U}$ and $\mathcal{V}$ as existing or non-existing edges. Here, the ground truth is that $\mathcal{U}$ contains positives and $\mathcal{V}$ the negatives. The nature of these can be homophilic or heterophilic, namely $\mathcal{U},\mathcal{V}\in\{hm,ht\}$. In Figure 4, we take into consideration Roman Empire as an example of a dataset where the models fail to achieve high performance, and we show that is not due to heterophilic edges (i.e. $AUC_{ht,hm},AUC_{ht,ht}$), since we have poor performance also on the homophilic ones $AUC_{hm,ht},AUC_{hm,hm}$. In Appendix -F, we provide additional evidence of this behavior, which also manifests when models reach high performance. Generally, we understand that the performance does not vary significantly, and the models’ rankings remain consistent. We conclude that models that achieve competitive performances in homophilic link prediction do not perform as well in our benchmark, but this is not due to class heterophily. This result answers negatively to RQ3 and highlights the need for new homophilic measures that can better identify datasets representative of link prediction under heterophily, because of its relevance in several applications.
@@ VI Conclusions
This paper presented GRAFF-LP, a link prediction model built upon GNNs and driven by a PIrd bias. Our method is designed to handle heterophilic graphs, enriching the current literature available on the topic, which is often confined to node classification. We show that GRAFF-LP outperforms the other baselines in 3 out of 4 datasets. We also propose $f_{g}$, a PIrd readout function based on the edge gradients, that, without any tuning, is able to improve the GRAFF-LP performance. Surprisingly, we observe that other models also benefit from $f_{g}$, but GRAFF-LP consistently ranks among the top 2. To better understand the effect of $f_{g}$, we study the evolution of the squared norm of the edge gradients via a novel metric named $GS$, and we surprisingly found that, under the lens of this norm, the GRAFF-LP’s message-passing separates the positive and negative edges. GRAFF-LP showcases this behavior more consistently than the other baselines, favored by the attraction and repulsion bias of the gradient flow interpretation. We also brought some evidence that the baselines do not struggle to perform link prediction because of class heterophily. This result exposes the need for new measures that help identify challenging scenarios to advance our knowledge of link prediction under heterophily. Future works should address the aspects related to new heterophily metrics based on the input features since they are not task-dependent. Moreover, we plan to further understand how GRAFF-LP can learn the edge gradient separation, even though it was not trained explicitly to do so. Then, it would be interesting to adapt other PIrd GNNs or develop new ones specifically suited for link prediction.